I have a 30-60-90 triangle one side opposite the hypotenuse is 3.000" long what are the other two sides? I thought 3.464 short side and hypotenuse 6.928 is this correct?
1
Expert's answer
2019-11-25T05:01:12-0500




No, not right.


c is the hypotenuse.


if a = 3.000, then

с = a/sin(30o) = 6.000

"b = \sqrt {c^2-a^2}"

b = 5.196


if b = 3.000, then

с = b/sin(60o) = 3.464

a = c*sin(30o) = 1.732